OpenStudy (anonymous):

7. A boat in calm seas travels in a straight line and ends the trip 22 km west and 53 km north of its original position. To the nearest tenth of a degree, find the direction of the trip.

OpenStudy (ivancsc1996):

Ok the tangent of the angle (theta) is \[\tan \theta=\frac{ a }{ b }=\frac{ 22km }{ 53km } \rightarrow \theta=\tan ^{-1}(\frac{ 22 }{ 53 })\]

OpenStudy (anonymous):

is this the whole problem?

OpenStudy (ivancsc1996):

 Yeah, I don't have a calculator so I can't give you the exact answer sorry but that's the whole problem.

OpenStudy (anonymous):

I still dont get this can someone help

OpenStudy (anonymous):

@radar

OpenStudy (anonymous):

@ivancsc1996

OpenStudy (radar):

Do you have a calculator?,